OpenStudy (ashleyisakitty):

Find the center and radius of the circle with equation (x+2)^2+(y+10)2=4. I am stuck, help?

OpenStudy (ashleyisakitty):

(x+2)^2+(y+10)^2=4 (x - (-2))^2 + (y - (10))^2 = 4 (x- (-2))^ + (y-10))^2 =???

OpenStudy (anonymous):

(x- (-2))^2 + (y- (-10))^2 = (2)^2 so if I'm not wrong results will be: C=(−2,−10) r=2

OpenStudy (anonymous):

annnas is right. Just look up equation of a circle

OpenStudy (ashleyisakitty):

wonderful, thanks guys

OpenStudy (anonymous):

Kittykitty you are smart :D

OpenStudy (anonymous):

(x-h)^2+(y-k)^2=r^2 centre is (h,k) and radius =r.

OpenStudy (anonymous):

O.O

OpenStudy (anonymous):

what surji said :)